# Hydrostatic Newton-Cartan Membranes

### v1

@@ IV Equilibrium Constituve Relations
An equilibrium state is specified by a given set of gauge invariant physical parameters $\{T,\mu,u^{\rho}\}$ identified with the temperature, mass chemical potential, and four velocity respectively. In equilibrium the associated generating function $\mathcal{F}[T,\mu,u]$ should be a function of the background geometry and the symmetry parameter set $\{V^{\mu},\Lambda^{V}\}$. This means the physical parameters have a solution in terms of the background geometry and symmetry set, such solution can easily be found to be Armas:2020 ; Jain:2020
$u^{a}=\frac{V^{a}}{V^{c}\tau_{c}}\,,\qquad T=\frac{T_{0}}{V^{c}\tau_{c}}\,,\qquad\frac{\mu}{T}=-\frac{1}{2}\frac{\Lambda^{V}}{T_{0}}+\frac{1}{2T}\bar{h}_{cd}u^{c}u^{d},$ (71)
where $T_{0}$ is a reference temperature, and $u^{\mu}$ has been normalised to $u^{\mu}\tau_{\mu}=1$. All we have to do now is write down the most general generating function $\mathcal{F}[T,\mu,K]$ up to an arbitrary number of derivatives. The equilibrium conditions (70) give rise to the conditions
@@
with $a^{d}$ as the acceleration, $\theta$ as the compresibility, $\sigma^{fg}$ as the shear, and $\Omega^{fg}$ as the vorticity. To finalize we can note the following definitions that will come in handy later
@@
$\bar{a}^{2}$ $\equiv\bar{h}_{cd}a^{c}a^{d}.$ (84)
IV.1 Equilibrium Partition Function
We need to classify all possible gauge invariant scalars that can be constructed using $\{\mu,T,u^{a},K_{ab},\bar{h}_{cd},\tau_{d}\}$ up to second order in derivatives [^7] The Canham-Helfrich functional used to derive the biconcave shape of blood cells contains terms of quadratic order in the extrinsic curvature meaning that for consistency we must keep all terms up to the same order in derivatives.. We are assuming both $\{T,\mu\}$ are zeroth order while $K_{ab}$ is taken to be first order.
@@ IV.1 Equilibrium Partition Function
$S_{(1)}=\left(K_{ab}h^{ab}\right)\left(K_{cd}h^{cd}\right)$ $S_{(2)}=\left(K_{ab}K_{cd}h^{ac}h^{bd}\right)$ $S_{(3)}=\left(K_{ab}K_{cd}u^{a}h^{bc}u^{d}\right)$ $S_{(4)}=\left(K_{ab}u^{a}u^{b}\right)\left(K_{cd}u^{c}u^{d}\right)$
$S_{(5)}=a^{c}a^{d}\bar{h}_{cd}$ $S_{(6)}=\left(K_{cd}u^{d}\right)a^{c}$ $S_{(7)}=\Omega^{ab}\Omega^{cd}\bar{h}_{ac}\bar{h}_{bd}$
Table 1: Summary of zeroth, first and second order independent inequivalent scalars which may contribute to the equilibrium partition function.
Where we assumed zero torsion and a flat ambient spacetime, ignore terms related to each other through partial integration, and used the corresponding integrability condition
@@
to ignore any explicit curvature terms. The equilibrium partition function can then be written as
$\mathcal{F}$ $=\int d^{d-1}\sigma\hat{e}\left[\chi(T,\mu)+\sum_{i=1}^{2}a^{(1)}_{i}(T,\mu)F_{(i)}+\sum_{i=1}^{7}a^{(2)}_{i}(T,\mu)S_{(i)}\right].$ (86)
Where $\chi(T,\mu)$ is the surface tension, and $\{a^{(1)}_{i}(T,\mu),a^{(2)}_{i}(T,\mu)\}$ are generic functions of the temperature and chemical potential. From this point onward, and for the sake of simplicity, we will omit the functional dependence of these aforementioned functions. The variation of the equilibrium partition function will take the form
@@
$+\int d^{d-1}\sigma\hat{e}\left[s\delta T+n\delta\mu+\sum_{i=1}^{2}\left(s^{(1)}_{i}\delta T+n^{(1)}_{i}\delta\mu\right)F_{(i)}+\sum_{i=1}^{7}\left(s_{i}^{(2)}\delta T+n_{i}^{(2)}\delta\mu\right)S_{(i)}\right]$
@@
where we define surface entropy density $s$, surface particle density $n$, and the surface susceptibilities $\{s^{(I)}_{i},n^{(I)}_{i}\}$
$s$ $=\left(\frac{\partial\chi}{\partial T}\right)_{\mu},\qquad$ $s^{(I)}_{i}$ $=\left(\frac{\partial a^{(I)}_{i}}{\partial T}\right)_{\mu},$ (88)
$n$ $=\left(\frac{\partial\chi}{\partial\mu}\right)_{T},\qquad$ $n^{(I)}_{i}$ $=\left(\frac{\partial a^{(I)}_{i}}{\partial\mu}\right)_{T},$
with $I=\{1,2\}$, and where we used the identity
$\delta\hat{e}$ $=\frac{\hat{e}}{2}\left(h^{ab}\delta\bar{h}_{ab}-2\upsilon^{a}\delta\tau_{a}\right).$ (89)
To find the relevant currents we will use the following variations of (71), (77), and (80)
@@
$\delta a^{d}$ $=\delta u^{c}D_{c}u^{d}+u^{c}\left(D_{c}\delta u^{d}+\delta\Gamma^{d}_{cf}u^{f}\right)=-\delta h^{cd}D_{c}\mu-h^{cd}D_{c}\delta\mu\,,$ (93)
$\delta\Omega^{fg}\Omega^{rs}\bar{h}_{fr}\bar{h}_{gs}$ $=\frac{1}{2}\left[\bar{u}_{a}\bar{h}_{bc}\Omega^{ab}-D_{c}\left(\mu+\frac{1}{2}\bar{u}^{2}\right)\right]\Omega^{cd}\delta\tau_{d}+\frac{1}{2}\left[\Omega^{af}\Omega^{bg}\bar{h}_{ab}\bar{h}_{fg}\right]u^{a}\delta\tau_{a}$ (94)
$-\frac{1}{2}\delta\bar{h}_{cd}\Omega^{cf}\Omega^{dg}\bar{h}_{fg}-\frac{1}{2}\bar{h}_{bc}\Omega^{cf}u^{b}u^{d}D_{f}\delta\tau_{d}-\frac{1}{2}\Omega^{cd}u^{b}D_{c}\delta h_{bd}\,,$
$\delta\Gamma^{a}_{cd}$ $=-\hat{\upsilon}^{a}D_{c}\delta\tau_{d}+\frac{1}{2}\delta h^{ab}\left(2D_{(c}\bar{h}_{d)b}-D_{b}\bar{h}_{cd}\right)+\frac{1}{2}h^{ab}\left(D_{(c}\delta\bar{h}_{d)b}-D_{b}\delta\bar{h}_{cd}\right),$ (95)
as well as the geometric identities
@@
Using (60) the surface energy current is found to be
$\left(\mathcal{T}_{\mathcal{S}}\right)^{c}$ $=-\chi\hat{\upsilon}^{c}-\left(\epsilon+\chi+\frac{n}{2}\bar{u}^{2}\right)u^{c}-\left(\sum_{i=1}^{2}\zeta^{(1)c}F_{(i)}+\sum_{i=1}^{7}\zeta^{(1)c}S_{(i)}\right)$ (98)
$+\left[2\left(a_{1}^{(1)}+2a_{1}^{(2)}F_{(1)}\right)\hat{\upsilon}^{a}K_{ab}+4a_{2}^{(2)}\hat{\upsilon}^{a}\left(K_{ar}h^{rs}K_{sb}\right)+2a_{3}^{(2)}\hat{\upsilon}^{a}\left(K_{ar}u^{r}\right)\left(K_{db}u^{d}\right)\right]h^{bc}$
$-\left[4a^{(2)}_{5}\hat{a}_{f}D_{g}\mu+2a^{(2)}_{6}K_{af}u^{a}D_{g}\mu\right]\upsilon^{(f}h^{g)c}$
@@
$+\left[2a^{(2)}_{(7)}\bar{u}_{a}\bar{h}_{bd}\Omega^{ab}-2a^{(2)}_{(7)}D_{d}\left(\mu+\frac{1}{2}\bar{u}^{2}\right)\right]\Omega^{cd}$
$+\left[2\left(a_{2}^{(1)}+2a_{4}^{(2)}F_{(2)}\right)K_{ab}u^{a}u^{b}-2a_{3}^{(2)}\left(K_{ar}h^{rs}K_{sb}\right)u^{a}u^{b}+a^{(2)}_{(6)}K_{fg}a^{f}u^{g}\right]u^{c}$
$+\left[D_{f}\left(a^{(2)}_{(7)}\bar{h}_{bc}\Omega^{cf}u^{b}\right)-a^{(2)}_{(7)}\Omega^{af}\Omega^{bg}\bar{h}_{ab}\bar{h}_{fg}\right]u^{c},$
the surface Cauchy stress-mass tensor is found to be
@@
$-\left[2\left(a_{1}^{(1)}+2a_{1}^{(2)}F_{(1)}\right)K_{ab}+4a_{2}^{(2)}\left(K_{ar}h^{rs}K_{sb}\right)+2a_{3}^{(2)}\left(K_{ar}u^{r}\right)\left(K_{bs}u^{s}\right)\right]h^{ac}h^{bd}$
$+\left[4a^{(2)}_{5}a_{f}D_{g}\mu+2a^{(2)}_{6}K_{af}u^{a}D_{g}\mu\right]h^{cf}h^{dg}+\left[a^{(2)}_{5}2h^{fg}D_{f}a_{g}+a^{(2)}_{6}h^{bf}D_{f}\left(K_{ab}u^{a}\right)\right]u^{c}u^{d}$
$+\left[2a^{(2)}_{7}\Omega^{ra}\Omega^{sb}\bar{h}_{rf}\bar{h}_{sg}\bar{h}_{ab}\right]h^{cf}h^{dg}-2D_{b}\left(a^{(2)}_{7}\Omega^{b(c}u^{d)}\right),$
and the bending moment is given by
@@
$\mathcal{Y}^{cdab}$ $=2a_{1}^{(2)}h^{cd}h^{ab}+2a_{2}^{(2)}h^{ac}h^{bd}+2a_{3}^{(2)}u^{(c}h^{d)b}u^{a}+2a_{4}^{(2)}u^{c}u^{d}u^{a}u^{b}.$ (101)
Where we have defined the internal energy $\epsilon$ as Armas:2020
$\epsilon=-\chi+Ts+n\mu,$ (102)
and the tensors $\{\zeta^{(I)c}_{i},\xi^{(I)cd}_{i}\}$ as
$\zeta^{(a)c}_{i}$ $=a^{(I)}_{i}\hat{\upsilon}^{c}+L^{(I)}_{i}u^{c}\,,$ (103)
$\xi^{(a)cd}_{i}$ $=a^{(I)}_{i}h^{cd}+n^{(I)}_{i}u^{c}u^{d}\,,$ (104)
with
$L^{(I)}_{i}$ $=s_{i}^{(I)}T+n_{i}^{(I)}\mu+\frac{1}{2}n_{i}^{(I)}\bar{u}^{2}.$ (105)
Using (88), the relation (102) is equivalent to the thermodynamics first law
$d\epsilon$ $=Tds+\mu dn$ (106)
IV.2 Axisymmetric configurations in 3+1 flat ambient spacetime
For applications concerning biological systems, such as the lipid bilayer membranes discussed in Helfrich ; Armas:2020 . It is enough to restrict the ambient spacetime geometry to
$\tau_{\mu}=\delta_{\mu}^{t}\,,\qquad h_{\mu\nu}$ $=\delta^{i}_{\mu}\delta_{\nu}^{j}\delta_{ij}\,,\qquad\hat{\upsilon}^{\rho}=-\delta^{\rho}_{t}\,,\qquad h^{\mu\nu}=\delta^{\mu}_{i}\delta^{\nu}_{j}\delta^{ij}\,,\quad m_{\mu}=0,$ (107)
where we are assuming a time direction $t$ and a Cartesian three-dimensional space with coordinates $x^{i}=(x,y,z)$. The equilibrium condition on this background, with gauge parameter $\Lambda^{K}=0$, has a solution for the Killing vector $V^{\mu}$ given by
$V=c_{0}\partial_{t}+c_{1}\left(z\partial_{y}-y\partial_{z}\right)+c_{2}\left(x\partial_{z}-z\partial_{x}\right)+c_{3}\left(y\partial_{x}-x\partial_{y}\right),$ (108)
where $\{c_{0},c_{1},c_{2},c_{3}\}$ are constants that can be chosen at our convenience. We will be interested in studying membranes subjected to axysimmetric flows, therefore we need to set $c_{1}=c_{2}=0$. This corresponds to a flow on the ambient spacetime with velocity

### v2

@@ 4 Equilibrium Constitutive Relations
An equilibrium state is specified by a given set of gauge invariant physical parameters $\{T,\mu,u^{\rho}\}$ identified with the temperature, the mass chemical potential, and the four velocity, respectively. In equilibrium, the associated generating function $\mathcal{F}[T,\mu,u]$ should be a function of the background geometry and the symmetry set $\{V^{\mu},\Lambda^{V}\}$. This means that the physical parameters have a solution in terms of the background geometry and the symmetry set; this solution can easily be found to be [3, 19]
$u^{a}=\frac{V^{a}}{V^{c}\tau_{c}}\,,\qquad T=\frac{T_{0}}{V^{c}\tau_{c}}\,$, (72)
$\frac{\mu}{T}=-\frac{1}{2}\frac{\Lambda^{V}}{T_{0}}+\frac{1}{2T}\bar{h}_{cd}u^{c}u^{d},$
where $T_{0}$ is a reference temperature, and $u^{\mu}$ has been normalized to $u^{\mu}\tau_{\mu}=1$. All that has to be done now is to write down the most general generating function $\mathcal{F}[T,\mu,K]$ up to an arbitrary number of derivatives. The equilibrium conditions (71) give rise to the conditions
@@
Where $a^{d}$ is the acceleration, $\theta$ the compresibility, $\sigma^{fg}$ the shear, and $\Omega^{fg}$ the vorticity. To finalize, we can note the following definitions that will come in handy later
@@
$\bar{a}^{2}$ $\equiv\bar{h}_{cd}a^{c}a^{d}=\bar{a}_{d}a^{d}\,$, (84)
$\bar{\Omega}^{2}$ $\equiv\Omega^{ab}\Omega^{cd}\bar{h}_{ac}\bar{h}_{bd}.$ (85)
4.1 Equilibrium Partition Function
We need to classify all possible gauge invariant scalars that can be constructed using $\{\mu,T,u^{a},K_{ab},\bar{h}_{cd},\tau_{d}\}$ up to second order in derivatives [^7] The Canham-Helfrich functional used to derive the biconcave shape of blood cells contains terms of quadratic order in the extrinsic curvature meaning that for consistency we must keep all terms up to the same order in derivatives.. We are assuming that both $\{T,\mu\}$ are zeroth order while $K_{ab}$ is taken to be first order.
@@ 4.1 Equilibrium Partition Function
$S_{(1)}=\left(K_{ab}h^{ab}\right)\left(K_{cd}h^{cd}\right)$ $S_{(2)}=\left(K_{ab}K_{cd}h^{ac}h^{bd}\right)$
$S_{(3)}=\left(K_{ab}K_{cd}u^{a}h^{bc}u^{d}\right)$ $S_{(4)}=\left(K_{ab}u^{a}u^{b}\right)\left(K_{cd}u^{c}u^{d}\right)$
$S_{(5)}=a^{c}a^{d}\bar{h}_{cd}$ $S_{(6)}=\left(K_{cd}u^{d}\right)a^{c}$
$S_{(7)}=\Omega^{ab}\Omega^{cd}\bar{h}_{ac}\bar{h}_{bd}$
Table 1: Summary of zeroth, first and second order independent inequivalent scalars which may contribute to the equilibrium partition function. Note that the scalars are independent before taking into account the transverse constraint (69). The equation of motion (69) will formally relate the scalars $F_{(1)}$ and $F_{(2)}$ with each other, leaving only one of them as independent. Consequently, $S_{(1)}$ and $S_{(4)}$ will also be related to each order.
When constructing table 1 we assumed zero torsion and a flat ambient spacetime. We also ignored terms related to each other through partial integration and used the corresponding integrability condition
@@
to ignore any explicit curvature terms. Note that the structures in table 1 have already been classified in [5, 4] on a relativistic setting. Namely, the scalars shown in table 1 can be regarded as the non-relativistic limit of those already obtained in [5, 4]. The equilibrium partition function can then be written as
$\mathcal{F}=\int d^{d-1}\sigma\hat{e}$ $\left[\chi(T,\mu)+\sum_{i=1}^{2}a^{(1)}_{i}(T,\mu)F_{(i)}+\right$. (87)
$\left.+\sum_{i=1}^{7}a^{(2)}_{i}(T,\mu)S_{(i)}\right].$
Where $\chi(T,\mu)$ is related to the surface tension  [^8] See equation (140) for the relation between $\chi$ and the surface tension and the spontaneous curvature of a biophysical membrane., and $\{a^{(1)}_{i}(T,\mu),a^{(2)}_{i}(T,\mu)\}$ are generic functions of the temperature and chemical potential. From this point onward, and for the sake of simplicity, we will omit the functional dependence of these aforementioned functions. Before moving forward, it is instructive to note that (87) can be written as
$\mathcal{F}$ $=\int d^{d-1}\sigma\hat{e}\left[\chi+a^{(1)}_{1}K+a^{(1)}_{2}\left(a\cdot n\right)+a^{(2)}_{1}K^{2}\right$. (88)
$a^{(2)}_{2}\left(K\cdot K\right)+a^{(2)}_{3}\left(n\cdot\Omega\right)^{2}+a^{(2)}_{4}\left(a\cdot n\right)^{2}+a^{(2)}_{5}\bar{a}^{2}$
$\left.\vphantom{\int}+a^{(2)}_{6}\left(\bar{a}\cdot\left(n\cdot\Omega\right)\right)+a^{(2)}_{7}\bar{\Omega}^{2}\right],$
where we used the following shorthand notation for the spatial traces of the extrinsic curvature
$K=h^{cd}K_{cd},\qquad\left(K\cdot K\right)=h^{ac}h^{bd}K_{ab}K_{cd},$ (89)
as well as the following shorthand notation for the normal components of acceleration and vorticity  [^9] The normal components of acceleration and vorticity show up due to the identities $u^{\mu}n_{\mu}=0$ and $K_{ab}u^{b}=n_{\mu}D_{a}u^{\mu}$. In equilibrium it holds that $h^{ac}K_{ab}u^{b}=n_{\rho}u^{c}_{\sigma}\Omega^{\rho\sigma}$ and $K_{ab}u^{a}u^{b}=n_{\mu}u^{\nu}\nabla_{\nu}u^{\mu}$
$n_{\nu}u^{\mu}\nabla_{\mu}u^{\nu}=n_{\nu}a^{\nu}$ $\equiv\left(a\cdot n\right),$ (90)
$n_{\rho}\Omega^{\rho\beta}$ $\equiv\left(n\cdot\Omega\right)^{\beta},$ (91)
and the following shorthand notation for some relevant contractions
$\left(n\cdot\Omega\right)^{\alpha}\bar{h}_{\alpha\beta}\left(n\cdot\Omega\right)^{\beta}$ $\equiv\left(n\cdot\Omega\right)^{2},$ (92)
$\left(n\cdot\Omega\right)^{\alpha}\bar{h}_{\alpha\beta}a^{\beta}$ $\equiv\left(\bar{a}\cdot\left(n\cdot\Omega\right)\right).$ (93)
From (88) we can note that the coefficients $\{a^{(1)}_{1},a^{(2)}_{1},a^{(2)}_{2}\}$ will be associated with the contributions of the extrinsic and intrinsic curvatures to the equilibrium partition function. Coefficients $\{a^{(1)}_{2},a^{(2)}_{4}\}$ will be associated with the contributions of the normal acceleration to the equilibrium partition function. The coefficient $a^{(2)}_{3}$ will be associated with the contribution of the normal vorticity to the equilibrium partition function. The coefficient $a^{(2)}_{5}$ will be associated with the contribution of the tangential acceleration to the equilibrium partition function. The coefficient $a^{(2)}_{7}$ will be associated with the contribution of the tangential vorticity to the equilibrium partition function. Finally, $a^{(2)}_{6}$ will be associated with the contribution of the spatial inner product between the normal vorticity and the tangential acceleration. After taking into account this form for $\mathcal{F}$, we note that the variation of the equilibrium partition function will take the form
@@
$+\int d^{d-1}\sigma\hat{e}\left[s\delta T+n\delta\mu+\sum_{i=1}^{2}\left(s^{(1)}_{i}\delta T+n^{(1)}_{i}\delta\mu\right)F_{(i)}\right$.
$\left.+\sum_{i=1}^{7}\left(s_{i}^{(2)}\delta T+n_{i}^{(2)}\delta\mu\right)S_{(i)}\right]$
@@
where we used the identity $\delta\hat{e}=\frac{\hat{e}}{2}\left(h^{ab}\delta\bar{h}_{ab}-2\upsilon^{a}\delta\tau_{a}\right)$, and defined the surface entropy density $s$, the surface particle density $n$, and the surface susceptibilities $\{s^{(1)}_{i},s^{(2)}_{i},n^{(1)}_{i},n^{(2)}_{i}\}$ as
$s$ $=\left(\frac{\partial\chi}{\partial T}\right)_{\mu}\,$$\qquad$ $n$ $=\left(\frac{\partial\chi}{\partial\mu}\right)_{T}\,$, (95)
$s^{(1)}_{i}$ $=\left(\frac{\partial a^{(1)}_{i}}{\partial T}\right)_{\mu}\,$$\qquad$ $n^{(1)}_{i}$ $=\left(\frac{\partial a^{(1)}_{i}}{\partial\mu}\right)_{T}\,$,
$s^{(2)}_{i}$ $=\left(\frac{\partial a^{(2)}_{i}}{\partial T}\right)_{\mu}\,$$\qquad$ $n^{(2)}_{i}$ $=\left(\frac{\partial a^{(2)}_{i}}{\partial\mu}\right)_{T}.$
To find the relevant currents we will use the following variations of (72)
@@
the variation of (78) given by
$\delta a^{d}$ $=\delta u^{c}D_{c}u^{d}+u^{c}\left(D_{c}\delta u^{d}+\delta\Gamma^{d}_{cf}u^{f}\right)$ (99)
$-\delta h^{cd}D_{c}\mu-h^{cd}D_{c}\delta\mu\,$,
also the variation of (81)
$\delta\Omega^{fg}\Omega^{rs}\bar{h}_{fr}\bar{h}_{gs}$ $=\frac{1}{2}\left[\Omega^{af}\Omega^{bg}\bar{h}_{ab}\bar{h}_{fg}\right]u^{a}\delta\tau_{a}$ (100)
$+\frac{1}{2}\left[\bar{u}_{a}\bar{h}_{bc}\Omega^{ab}-D_{c}\left(\mu+\frac{1}{2}\bar{u}^{2}\right)\right]\Omega^{cd}\delta\tau_{d}$
$-\frac{1}{2}\delta\bar{h}_{cd}\Omega^{cf}\Omega^{dg}\bar{h}_{fg}-\frac{1}{2}\bar{h}_{bc}\Omega^{cf}u^{b}u^{d}D_{f}\delta\tau_{d}$
$-\frac{1}{2}\Omega^{cd}u^{b}D_{c}\delta h_{bd}\,$,
as well as the variation of the connection
$\delta\Gamma^{a}_{cd}$ $=-\hat{\upsilon}^{a}D_{c}\delta\tau_{d}+\frac{1}{2}\delta h^{ab}\left(2D_{(c}\bar{h}_{d)b}-D_{b}\bar{h}_{cd}\right)$ (101)
$+\frac{1}{2}h^{ab}\left(D_{(c}\delta\bar{h}_{d)b}-D_{b}\delta\bar{h}_{cd}\right),$
and the geometric identities
@@
Using (60), the surface energy current is found to be
$\left(\mathcal{T}_{\mathcal{S}}\right)^{c}$ $=-\chi\hat{\upsilon}^{c}-\left(\epsilon+\chi+\frac{n}{2}\bar{u}^{2}\right)u^{c}-\left(\sum_{i=1}^{2}\zeta^{(1)c}_{i}F_{(i)}+\sum_{i=1}^{7}\zeta^{(2)c}_{i}S_{(i)}\right)$ (104)
$+\left[2\left(a_{1}^{(1)}+2a_{1}^{(2)}F_{(1)}\right)\hat{\upsilon}^{a}K_{ab}+4a_{2}^{(2)}\hat{\upsilon}^{a}\left(K_{ar}h^{rs}K_{sb}\right)\right$.
$\left.+2a_{3}^{(2)}\hat{\upsilon}^{a}\left(K_{ar}u^{r}\right)\left(K_{db}u^{d}\right)\right]h^{bc}$
$-\left[4a^{(2)}_{5}\bar{a}_{f}D_{g}\mu+2a^{(2)}_{6}K_{af}u^{a}D_{g}\mu\right]\upsilon^{(f}h^{g)c}$
@@
$+\left[2a^{(2)}_{7}\bar{u}_{a}\bar{h}_{bd}\Omega^{ab}-2a^{(2)}_{7}D_{d}\left(\mu+\frac{1}{2}\bar{u}^{2}\right)\right]\Omega^{cd}$
$+\left[D_{f}\left(a^{(2)}_{7}\bar{h}_{bc}\Omega^{cf}u^{b}\right)-a^{(2)}_{7}\Omega^{af}\Omega^{bg}\bar{h}_{ab}\bar{h}_{fg}\right]u^{c}$
$+\left[2\left(a_{2}^{(1)}+2a_{4}^{(2)}F_{(2)}\right)K_{ab}u^{a}u^{b}-2a_{3}^{(2)}\left(K_{ar}h^{rs}K_{sb}\right)u^{a}u^{b}\right$.
$\left.+a^{(2)}_{6}K_{fg}a^{f}u^{g}\right]u^{c},$
where the tensors $\{\zeta^{(1)c}_{i},\zeta^{(2)c}_{i}\}$ are defined as
$\zeta^{(1)c}_{i}$ $=a^{(1)}_{i}\hat{\upsilon}^{c}+\left[s^{(1)}_{i}T+n^{(1)}_{i}\left(\mu+\frac{1}{2}\bar{u}^{2}\right)\right]u^{c}\,$, (105)
$\zeta^{(2)c}_{i}$ $=a^{(2)}_{i}\hat{\upsilon}^{c}+\left[s^{(2)}_{i}T+n^{(2)}_{i}\left(\mu+\frac{1}{2}\bar{u}^{2}\right)\right]u^{c}\,$, (106)
and where we have defined the internal energy $\epsilon$ as [3]
$\epsilon=-\chi+Ts+n\mu,$ (107)
Note that using (95), the relation (107) is equivalent to the thermodynamics first law
$d\epsilon$ $=Tds+\mu dn.$ (108)
Using (60), the surface Cauchy stress-mass tensor is found to be
@@
$-\left[2\left(a_{1}^{(1)}+2a_{1}^{(2)}F_{(1)}\right)K_{ab}+4a_{2}^{(2)}\left(K_{ar}h^{rs}K_{sb}\right)+2a_{3}^{(2)}\left(K_{ar}u^{r}\right)\left(K_{bs}u^{s}\right)\right]h^{a(c}h^{d)b}$
$+\left[4a^{(2)}_{5}\bar{a}_{f}D_{g}\mu+2a^{(2)}_{6}K_{af}u^{a}D_{g}\mu\right]h^{cf}h^{dg}+\left[a^{(2)}_{5}2h^{fg}D_{f}\bar{a}_{g}+a^{(2)}_{6}h^{bf}D_{f}\left(K_{ab}u^{a}\right)\right]u^{c}u^{d}$
$+\left[2a^{(2)}_{7}\Omega^{ra}\Omega^{sb}\bar{h}_{rf}\bar{h}_{sg}\bar{h}_{ab}\right]h^{f(c}h^{d)g}-2D_{b}\left(a^{(2)}_{7}\Omega^{b(c}u^{d)}\right),$
where the tensors $\{\xi^{(1)cd}_{i},\xi^{(2)cd}_{i}\}$ are defined as
$\xi^{(1)cd}_{i}$ $=a^{(1)}_{i}h^{cd}+n^{(1)}_{i}u^{c}u^{d}\,$, (110)
$\xi^{(2)cd}_{i}$ $=a^{(2)}_{i}h^{cd}+n^{(2)}_{i}u^{c}u^{d}\,.$ (111)
Finally, using (60), the bending moment is found to be
@@
$\mathcal{Y}^{cdab}$ $=2a_{1}^{(2)}h^{cd}h^{ab}+2a_{2}^{(2)}h^{ac}h^{bd}$ (113)
$+2a_{3}^{(2)}u^{(c}h^{d)b}u^{a}+2a_{4}^{(2)}u^{c}u^{d}u^{a}u^{b}.$
The explicit form of the variations that were used to obtain (104),(109), and (112) are shown in appendix A. In this appendix, the individual contributions of each term to the currents is also shown.
4.2 Axisymmetric configurations in 3+1 flat ambient spacetime
For applications concerning biological systems, such as the lipid bilayer membranes discussed in [17, 3]. It is enough to restrict the ambient spacetime geometry to
$\tau_{\mu}$ $=\delta_{\mu}^{t}\,$$\qquad$ $h_{\mu\nu}$ $=\delta^{i}_{\mu}\delta_{\nu}^{j}\delta_{ij}\,$, (114)
$\hat{\upsilon}^{\rho}$ $=-\delta^{\rho}_{t}\,$, $h^{\mu\nu}$ $=\delta^{\mu}_{i}\delta^{\nu}_{j}\delta^{ij}\,$$\quad$ $m_{\mu}$ $=0,$
where we assume a time direction $t$ and a Cartesian three-dimensional space with coordinates $x^{i}=(x,y,z)$. The equilibrium condition on this background, with gauge parameter $\Lambda^{K}=0$, has a solution for the Killing vector $V^{\mu}$ given by
$V$ $=c_{0}\partial_{t}+c_{1}\left(z\partial_{y}-y\partial_{z}\right)+c_{2}\left(x\partial_{z}-z\partial_{x}\right)$ (115)
$+c_{3}\left(y\partial_{x}-x\partial_{y}\right),$
where $\{c_{0},c_{1},c_{2},c_{3}\}$ are constants that can be chosen at our convenience. We will be interested in studying membranes subjected to axisymmetric flows, therefore, we need to set $c_{1}=c_{2}=0$. This corresponds to a flow on the ambient spacetime with velocity